OpenStudy (anonymous):

how do you find the height of a right triangle if you only know the base and 1 angle using tan, cos, our sin ?

OpenStudy (australopithecus):

just use SOH CAH TOA sin = opposite/Hypotenuse cos = adjacent/hypotenuse tan = opposite/adjacent if you have the angle and the base you can just rearrange one of these formulas to solve for the height for example 

OpenStudy (australopithecus):

We have the adjacent = 23 and the opposite = x the angle is 32 degrees thus we use tan(x) = opposite/adjacent thus we have tan(32) =23/x xtan(32) = 23 x = 23/tan(32)

OpenStudy (australopithecus):

x = 36.8

OpenStudy (anonymous):

Ohhh that makes since!!

OpenStudy (australopithecus):

furthermore if you know 2 angles in the triangle you can easily figure out the third remember all the angles of a triangle add up to 180